OpenStudy (anonymous):

Find the product, if possible. AB, if A= [0 -5 10] B= [-9] [-1 6 -9] [0] , [4]

OpenStudy (amistre64):

0.-9 + -5.0 + 10.4 = 40 9-36 = -27 [ 40] [-27] i think

OpenStudy (anonymous):

I am not certain about what I'm looking at. if A= [0 -5 10] B= [-9] What does the rest mean? Doing matrix multiplication is easy enough, but the form in which it is written is a bit confusing.